OpenStudy (anonymous):

It is important that physicians know how long a medicine will stay in a person’s body to keep a sufficient amount but also so that the person does not overdose. For example, a patient takes a pill that has 50 mg (milligrams) dosage and wakes up the next day with a particular percentage of the medication washed out of his or her system.

myininaya (myininaya):

Is there a question here? A mathematical question?

OpenStudy (anonymous):

Demonstrate your knowledge of conversions by first choosing a percentage of the medicine that is washed out the next day, and convert this percentage to a decimal and then a fraction. How much medication remains in the body after the percentage amount of your choosing is washed out? Explain which conversion (percent, decimal, or fraction) was the easiest to use to make the calculation. Show all your calculations. Real Life

OpenStudy (anonymous):

Thats the second part

myininaya (myininaya):

so we get to choose the percentage that is washed out. Ok how about we choose it to be 50%. 50%=.5=1/2

myininaya (myininaya):

why do we get to choose how much is washed out?

myininaya (myininaya):

that is weird to me

OpenStudy (anonymous):

I do not know but im so sorry i put up the wrong question check the other one please

OpenStudy (anonymous):

Robert is interested in buying carpet for his house. Robert measured the areas where he plans to install the carpet and found that the hallway needed a total of 60 square feet, the living room needed 250 square feet, and the master bedroom required 120 square feet. However, when Robert called to order the carpet, he found out that he had to order in square yards. Research to find the conversion between square feet and square yards to answer the following: How many square yards total will Robert need to buy to carpet his house? What percentage of the total carpet will be used in the master bedroom? Convert the percent obtained for the master bedroom into a decimal and then to a reduced fraction. Show all of your calculations.

OpenStudy (anonymous):

??? which question are you doing?

OpenStudy (anonymous):

The one that starts with robert

OpenStudy (anonymous):

how many square feet total? 60 + 250 + 120 = 430 right?

OpenStudy (anonymous):

Yes but you have to answer the questions underneath it

OpenStudy (anonymous):

430 square feet, divide by 9 get \[47\tfrac{7}{9}\] yards

OpenStudy (anonymous):

fraction of total used in bedroom is \[\frac{250}{430}=\frac{25}{43}\]

OpenStudy (anonymous):

this does not reduce further. you sure these numbers are right? because the question is goofy

OpenStudy (anonymous):

Yes well are all those measured in yards

OpenStudy (anonymous):

no that was square feet.

OpenStudy (anonymous):

but as a proportion it makes no difference which you use

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Hey one more question how would make the bedroom answer into a percent

OpenStudy (king):

25/43 * 100/430?